# Bidirectional Gated Mamba for Sequential Recommendation

### v2

@@ 2.5. Train and Inference
3. Experiment
In this section, first we introduce the experiment setting. Then we present extensive experiments to evaluate the effectiveness of our proposed SIGMA. All reported experimental results are the average values obtained from five independent runs of the algorithm
3.1. Experiment Setting
3.1.1. Dataset
We conduct comprehensive experiments on five representative real-world datasets i.e., Yelp [^1] https://www.yelp.com/dataset, Amazon-based [^2] https://cseweb.ucsd.edu/jmcauley/datasets.html#amazon_reviews (Beauty, Sports and Games) and MovieLens-1M [^3] https://grouplens.org/datasets/movielens/. The statistics of datasets after preprocessing are shown in Table 1. For the grouped user analysis, all datasets are categorized into three subsets based on user interaction length: “Short” ($0-5$), “Medium” ($5-20$), and “Long” ($20-inf$). Additionally, we arrange user interactions sequentially by time across all datasets and use the leave-one-out method for data splitting. The detailed description and preprocessing procedures can be found in Appendix B.
3.1.2. Evaluation Metrics
To assess performance, we use Top-10 Hit Rate (HR@10), Top-10 Normalized Discounted Cumulative Gain (NDCG@10), and Top-10 Mean Reciprocal Rank (MRR@10) as evaluation metrics, all of which are widely used in related studies (Gu and Dao, 2023; Jiang et al., 2024; De et al., 2024). These metrics offer a comprehensive evaluation of the SRS’s performance. For more details, please refer to the description and calculation methods in Appdendix C.
3.1.3. Implementation Details
In this section, we provide a detailed description of our framework’s implementation. The code can be found in Supplementary Material. For GPU selection, all experiments are conducted on a single NVIDIA L4 GPU. The Adam optimizer (Kingma and Ba, 2014) is used with a learning rate of 0.001. For a fair comparison, the embedding dimension for all tested models is set to 64. Other implementation details are the same as original papers (Liu et al., 2024a; Wang et al., 2024; Kang and McAuley, 2018), which are presented in Appendix D.
3.1.4. Baselines
To demonstrate the effectiveness and efficiency of our proposed framework, we compare SIGMA with state-of-the-art transformer-based models (BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), LinRec (Liu et al., 2023a), FEARec (Du et al., 2023)), RNN-based models (GRU4Rec (Jannach and Ludewig, 2017)), and SSM-based models (Mamba4Rec (Liu et al., 2024a), denoted as Mamba, ECHOMamba4Rec (Wang et al., 2024)), denoted as ECHO. Additional details are provided in Appendix E.
Table 3. Efficiency comparison of inference time per batch (ms) and GPU memory usage (GB).
@@ 3.1.4. Baselines
3.2. Overall Performance Comparison
@@
From the results, RNN-based models struggle with complex dependencies, resulting in relatively inferior performance. Besides, transformer-based models often show comparable performance, suggesting their powerful capacities in sequence modeling by self-attention. However, they still slightly lag behind our SIGMA because of the short sequence modeling problem they are facing and Mamba’s more powerful abilities in capturing long-term dependency (Yang et al., 2024). In terms of the SSM-based models, we find that they also underperform our SIGMA consistently, because of the context modeling and short sequence modeling problems mentioned before. Especially, Mamba4Rec and ECHOMamba4Rec show inferior performance in the Sports and Beauty datasets, whose average lengths are relatively shorter. Such a phenomenon emphasizes their weaknesses in long-tail users by direct adaptation of Mamba for the sequential recommendation.
3.3. Efficiency Comparison
@@
In this section, we analyze the efficiency of SIGMA compared to other baselines by examining the inference time per batch and GPU memory usage during inference. The results, presented in Table 3, offer several valuable insights. First, we can find that the Mamba-based methods, including our SIGMA, can achieve higher efficiency remarkably compared with the transformer-based methods, except for LinRec. The reason lies in the simple input-dependent selection mechanism of Mamba. Then, though the efficiency-specified LinRec also owns comparable efficiency, it slightly downgrades the effectiveness of the transformer. By comparison, our SIGMA can achieve a better efficiency-effectiveness trade-off. To further analyze the efficiency of our SIGMA from a theoretical view, we refer to the complexity analysis in Appendix A. Besides, Additional experiments on the other three datasets are provided in Appendix F.
Figure 3. User group analysis on Beauty and Sports.
3.4. Grouped Users Analysis
@@
This section presents the recommendation quality for users with different number of interactions to figure out the effctiveness of SIGMA on enhancing long-tail users’ experience. We illustrate the results on Beauty and Sports in Figure 3. More results on other datasets are in Appendix G. From this figure, we can find that:
@@
  • Mamba4Rec (denoted as Mamba in the figure) that adopts the vanilla Mamba structures for SRS presents the poor performance for ‘short’ and ‘medium’ users. While ECHO, which designs a bi-directional modeling module for SRS, achieves a slightly better results while is still worse than SASRec, a SRS baseline.
@@
  • Our SIGMA defeat all baselines on all groups, where FE-GRU contributes to the short-sequence modeling and PF-Mamba boosts the overall performance.
3.5. Ablation Study
@@
In this section, we analyze the efficacy of three key components within SIGMA, including PF-Mamba (partial flipping and DS gate), and FE-GRU. We design three variants: (1) w/o partial flipping: this variant uses the original interaction sequence without partial flipping; (2) w/o DS gate: the second variant linearly combines the output of two Mamba blocks; (3) w/o FE-GRU: this variant drops the Feature Extract GRU. We test these variants on Beauty and present results in Table 4 and Figure 4. Results on other datasets are included in Appendix H.
From Table 4 and Figure 4, we can conclude that:
@@
  • With the bi-directional interaction sequences, partial flipping contributes to improve the recommendation performance for all users.
@@
Figure 4. Ablation analysis on Beauty.
Table 4. Ablation study on Beauty.
@@
3.6. Hyperparameter Analysis
@@
Figure 5. Parameter study for $r$ on Beauty.
Table 5. Parameter study for $\mathit{L}$ on Beauty.
@@
3.7. Case Study
In this section, we leverage a specific example in ML-1M to illustrate the effectiveness of partial flipping in SIGMA. Specifically, we choose a user (ID: 5050) and present the interaction sequence before and after the partial flipping in the left part of Figure 6. With $r=1$, only the last item $2762$ is remained at the original position and other items are flipped. From this example, we can find that this user prefers comedy and romance movies (pink balls), as well as action and thriller movies (blue balls). Without the flipping, baselines focus on the most recent interactions on action and thriller movies and provide incorrect recommendations of the same genres (movie 3753 and 2028). While our SIGMA, with PF-Mamba, notices the previous preference on comedy and romance movies, makes the accurate recommendation of movie 539. Furthermore, we also present the overall performance for user-5050 in Table 6, where SIGMA significantly defeat baselines.
Figure 6. Case study for User-$5050$ in ML-1M.
Table 6. Performance comparison on User-$5050$.
@@ 3.7. Case Study
4. Related Work
4.1. Sequential Recommendation
Advancements in deep learning have transformed recommendation systems, making them more personalized and accurate in next-item prediction (Liu et al., 2023c, b, 2024b; Li et al., 2022b; Gao et al., 2024; Li et al., 2023a, b; Chen et al., 2023; Cui et al., 2024). Early sequential recommendation frameworks have adopted CNNs and RNNs to capture users’ preferences but faced issues like catastrophic forgetting when dealing with long-term dependencies (de Souza Pereira Moreira et al., 2021; Kim et al., 2019a). Then, the transformer-based models have emerged as powerful methods with their self-attention mechanism, significantly improving performance by selectively capturing the complex user-item interactions (Kang and McAuley, 2018). However, they have suffered from inefficiency due to the quadratic computational complexity (Keles et al., 2023). Therefore, to address the trade-off between effectiveness and efficiency, we propose SIGMA, a novel framework that achieves remarkable performance.
4.2. Selective State Space Model
Currently, SSM-based models have been proven effective in time-series prediction due to their ability to capture the hidden dynamics (Hamilton, 1994). To further address the issues of catastrophic forgetting and long-term dependency in sequential processing, a special SSM called Mamba was introduced. Originating from the structured state space model (S4)(Smith et al., 2022; Hamilton, 1994), Mamba has been proven to provide Transformer-level performance with better efficiency, particularly for long sequences (Yang et al., 2024). However, it still suffers from some challenges when adopted in the realm of recommendation i.e., context modeling and short sequence modeling, which are mainly caused by Mamba’s original structure and the inflexibility in hidden state transferring. Correspondingly, we introduce a special bi-directional module called Partially Flipped Mamba and a Feature Extract GRU in our SIGMA framework, which somewhat addresses these problems and explores a novel way to leverage Mamba in SRS.
5. Conclusion

### v3

@@ Introduction
Table 2: Overall performance comparison between SIGMA and other baselines. The best results are bold, and the second-best are underlined. “*” indicates the improvements are statistically significant (i.e., one-sided t-test with $p<0.05$ ) over baselines.
Experiment
In this section, we first introduce the experiment setting. Then, we present extensive experiments to evaluate the effectiveness of SIGMA.
Experiment Setting
Dataset. We conduct comprehensive experiments on five representative real-world datasets i.e., Yelp [^1] https://www.yelp.com/dataset, Amazon series [^2] https://cseweb.ucsd.edu/jmcauley/datasets.html“#amazon˙reviews (Beauty, Sports and Games) and MovieLens-1M [^3] https://grouplens.org/datasets/movielens/.The statistics of datasets after preprocessing are shown in Table 1. For the grouped user analysis, all datasets are categorized into three subsets based on user interaction length: “Short” ($0-5$), “Medium” ($5-20$), and “Long” ($20+$). Additionally, we arrange user interactions sequentially by time across all datasets.
Evaluation Metrics. To assess performance, we use Top-10 Hit Rate (HR@10), Top-10 Normalized Discounted Cumulative Gain (NDCG@10), and Top-10 Mean Reciprocal Rank (MRR@10) as evaluation metrics, all of which are widely used in related studies (Gu and Dao 2023; Jiang, Han, and Mesgarani 2024; De et al. 2024). These metrics offer a comprehensive evaluation of the SRS’s performance. All experimental results reported are averages from five independent runs of the framework.
Implementation Details. In this section, we provide a detailed description of our framework’s implementation. For GPU selection, all experiments are conducted on a single NVIDIA L4 GPU. The Adam optimizer (Kingma and Ba 2014) is used with a learning rate of 0.001. For a fair comparison, the embedding dimension for all tested models is set to 64. Other implementation details are the same as original papers (Liu et al. 2024a; Wang, He, and Zhu 2024; Kang and McAuley 2018).
Baselines. To demonstrate the effectiveness and efficiency of our proposed framework, we compare SIGMA with state-of-the-art transformer-based models (BERT4Rec (Sun et al. 2019), SASRec (Kang and McAuley 2018), LinRec (Liu et al. 2023a), FEARec (Du et al. 2023)), RNN-based models (GRU4Rec (Jannach and Ludewig 2017)), and SSM-based models (Mamba4Rec (Liu et al. 2024a), denoted as Mamba, ECHOMamba4Rec (Wang, He, and Zhu 2024), denoted as ECHO).
@@
Yelp SASRec 443ms 9.28G
FEARec 483ms 10.01G
LinRec 353ms 7.46G
Mamba 361ms 7.32G
ECHO 368ms 8.46G
SIGMA 352ms 8.27G
Table 3: Efficiency comparison of inference time per batch (ms) and GPU memory usage (GB).
Overall Performance Comparison
@@
From the results, RNN-based models struggle with complex dependencies, resulting in relatively inferior performance. Besides, transformer-based models often show comparable performance, suggesting their powerful capacities in sequence modeling by self-attention. However, they still slightly lag behind our SIGMA because of the short sequence modeling problem they are facing and Mamba’s more powerful abilities in capturing long-term dependency (Yang et al. 2024).
In terms of the SSM-based models, we find that they also underperform our SIGMA consistently, because of the context modeling and short sequence modeling problems mentioned before. Specifically, Mamba4Rec and ECHOMamba4Rec show inferior performance in the Sports and Beauty datasets, whose average lengths are relatively shorter. Such a phenomenon emphasizes their weaknesses in long-tail users by direct adaptation of Mamba for the sequential recommendation.
Efficiency Comparison
@@
In this section, we analyze the efficiency of SIGMA compared to other baselines by examining the inference time per batch and GPU memory usage during inference. The results, presented in Table 3, offer several valuable insights. First, we can find that the Mamba-based methods, including our SIGMA, can achieve higher efficiency remarkably compared with the transformer-based methods, except for LinRec. The reason lies in the simple input-dependent selection mechanism of Mamba. Then, though the efficiency-specified LinRec also owns comparable efficiency, it slightly downgrades the effectiveness of the transformer. By comparison, our SIGMA can achieve a better efficiency-effectiveness trade-off.
Figure 3: User group analysis on Beauty and Sports.
Grouped Users Analysis
@@
This section presents the recommendation quality for users with varying lengths of interaction histories, aiming to provide a deeper insight into SIGMA’s effectiveness in enhancing the experience of long-tail users. We illustrate the results on Beauty and Sports in Figure 3 and find that:
@@
  • Mamba4Rec that adopts the vanilla Mamba structures for SRS presents poor performance for ‘short’ and ‘medium’ users. While ECHO, which designs a bi-directional modeling module for SRS, achieves slightly better results while is still worse than SASRec.
@@
  • Our SIGMA defeats all baselines on all groups, where FE-GRU contributes to the short-sequence modeling and PF-Mamba boosts the overall performance.
Ablation Study
@@
In this section, we analyze the efficacy of three key components within SIGMA, including PF-Mamba (partial flipping and DS gate), and FE-GRU. We design three variants: (1) w/o partial flipping: this variant uses the original interaction sequence without partial flipping; (2) w/o DS gate: the second variant linearly combines the output of two Mamba blocks; (3) w/o FE-GRU: this variant drops the Feature Extract GRU. We test these variants on Beauty and present results in Table 4 and Figure 4. We can conclude that:
@@
  • With the bi-directional interaction sequences, partial flipping contributes to improving the recommendation performance for all users.
@@
Figure 4: Ablation analysis on Beauty.
@@
Table 4: Ablation study on Beauty.
Hyperparameter Analysis
@@
Figure 5: Parameter study for $r$ on Beauty.
@@
Case Study
In this section, we leverage a specific example in ML-1M to illustrate the effectiveness of partial flipping in SIGMA. Specifically, we choose a user (ID: 5050) and present the interaction sequence before and after the partial flipping in the left part of Figure 6. With $r=1$, only the last item $2762$ remained at the original position, and other items are flipped. From this example, we can find that this user prefers comedy and romance movies (pink balls), as well as action and thriller movies (blue balls). Without the flipping, baselines focus on the most recent interactions on action and thriller movies and provide incorrect recommendations of the same genres (movie 3753 and 2028). While our SIGMA, with PF-Mamba, notices the previous preference for comedy and romance movies, makes the accurate recommendation of movie 539. Furthermore, we also present the overall performance for user-5050 in Table 6, where SIGMA significantly defeats baselines.
Figure 6: Case study for User-$5050$ in ML-1M.
@@
Table 6: Performance comparison on User-$5050$.
Related Work
Sequential Recommendation
Advancements in deep learning have transformed recommendation systems, making them more personalized and accurate in next-item prediction (Liu et al. 2023b; Wang et al. 2024; Liu et al. 2024c). Early sequential recommendation frameworks have adopted CNNs and RNNs to capture users’ preferences but faced issues like catastrophic forgetting when dealing with long-term dependencies (de Souza Pereira Moreira et al. 2021; Kim et al. 2019a). Then, the transformer-based models have emerged as powerful methods with their self-attention mechanism, significantly improving performance by selectively capturing the complex user-item interactions (Li et al. 2023a). However, they have suffered from inefficiency due to the quadratic computational complexity (Keles, Wijewardena, and Hegde 2023). Therefore, to address the trade-off between effectiveness and efficiency, we propose SIGMA, a novel framework that achieves remarkable performance.
Selective State Space Model
Currently, SSM-based models have been proven effective in time-series prediction due to their ability to capture the hidden dynamics (Smith, Warrington, and Linderman 2022; Hamilton 1994). To further address the issues of catastrophic forgetting and long-term dependency in sequential processing, a special SSM called Mamba was introduced. Attributing to its unique selectivity (Gu and Dao 2023), Mamba shows remarkable performance without leveraging any sequence denoising methods (Zhang et al. 2023, 2022; Lin et al. 2023) or feature selecting methods (Lin et al. 2022), even when addressing long sequences (Yang et al. 2024). However, it still suffers from some challenges when adopted in the realm of recommendation i.e., context modeling and short sequence modeling, which are mainly caused by Mamba’s original structure and the inflexibility in hidden state transferring. Correspondingly, we introduce a special bi-directional module called Partially Flipped Mamba and a Feature Extract GRU in our SIGMA framework, which somewhat address these problems and explores a novel way to leverage Mamba in SRS.
Conclusion
@@
Acknowledgements
This research was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund No. 2024002), CCF-BaiChuan-Ebtech Foundation Model Fund, and Kuaishou.
Appendix
A. Complexity Analysis
In this section, we will analyze the time complexity of our proposed SIGMA framework. We denote training iteration as ${I}$, the sequence length as ${N}$, the embedding size as ${D}$ and the kernel size of Conv1d as $k$. Following the previous formulas, we analyze the time complexity of the key components: (i) Firstly, since our DS Gate consists of several linear layers and one Conv1d Layer, the computation cost of it can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times k\right)$. (ii) Secondly, the time complexity of Mamba is known as $\mathcal{O}\left(N\times D\right)$ (Gu and Dao 2023). In the PF-Mamba, we add one more reverse direction and several linear layers, therefore, the total complexity can be calculated as $\mathcal{O}\left(3\times N\times D^{2}+N\times D\times(k+2)\right)$. (iii) Thirdly, for the FE-GRU, which consists of Conv1d and GRU cell, the computational complexity can be calculated by simply adding GRU and Conv1d together: $\mathcal{O}\left(N\times D^{2}+N\times D\times k\right)$. In conclusion, the whole time complexity of SIGMA is $\mathcal{O}\left(13\times N\times D^{2}+N\times D\times(2k+2)\right)$. Considering the extreme situation when dealing with quite long sequences ($\left(N\gg D\right)$), it can be further simplified to $\mathcal{O}\left(N\right)$, compared to the $\mathcal{O}\left({N}^{2}\right)$ complexity of transformer-based models (Liu et al. 2024a), showing its superiority in efficiency, which also support by the experimental results on ML-1M dataset.
B. Dataset Information
We mainly evaluate our framework on five real-world datasets, i.e., Beauty, Sports, Games, Yelp and ML-1M, which are all large-scale public datasets that have been widely used as benchmarks in the next-item prediction task (Yang et al. 2024).
  • Yelp [^4] https://www.yelp.com/dataset: This dataset is released by Yelp as part of their Dataset Challenge. The data source includes user reviews, business information, and user interactions on Yelp. It contains over $6.9$ million reviews, details on more than $150,000$ businesses, and user interaction data like check-ins and tips.
  • Amazon based [^5] https://cseweb.ucsd.edu/jmcauley/datasets.html“#amazon˙reviews: These three datasets, provided by Amazon, include customer reviews and metadata from the Beauty, Sports, and Games categories which feature millions of reviews, with attributes such as review text, ratings, product IDs, and reviewer information.
  • MovieLens-1M [^6] https://grouplens.org/datasets/movielens/: The MovieLens 1M dataset is released by GroupLens Research. It comprises 1 million movie ratings from $6,041$ users on $3417$ movies with attributes like user demographics, movie titles, genres, and timestamps.
In the experiments, we employ a leave-one-out method for splitting the datasets and arrange all the user interactions sequentially by time. Moreover, for each user and item, we construct an interaction sequence by simply sorting their interaction records based on timestamps and ratings. Considering the average length and total samples of each dataset vary, we filter users and items with less than five recorded interactions for ML-1M, Beauty, and Games, following the setting in original papers (Qin, Yang, and Zhong 2024). For Yelp and Sports, we also added upper bounds ($100$ for Yelp and $200$ for Sports).
C. Evaluation Metrics
In this section, we will detail the information and calculations of our selected evaluation metrics.
  • HR@10: HR@10 represents Hit Rate truncated at 10, which measures the fraction of users for whom the correct item is ranked within the top 10 predictions. Specifically, for a user $u$, it is calculated as:
[…]
    where $v_{u}$ is the correct item for user $u$ and $\mathbb{1}(\cdot)$ is an indicator function that equals 1 if the condition is true and 0 otherwise.
  • NDCG@10: NDCG@10 represents Normalized Discounted Cumulative Gain truncated at 10 which evaluates the ranking quality by giving higher importance to items ranked at the top. It is defined as:
    IDCG@10 $=\sum_{k=1}^{10}\frac{2^{\text{rel}^{\ast}(v_{uk})}-1}{\log_{2}(k+1)}$ (14)
    NDCG@10 $=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\frac{1}{\text{IDCG@10}(v_{u})}\sum_{k=1}^{10}\frac{2^{\text{rel}(v_{uk})}-1}{\log_{2}(k+1)}$
[…]
  • MRR@10: MRR@10 represents Mean Reciprocal Rank truncated at 10, which measures the average rank position of the first relevant item. For user $u$, it is defined as:
[…]
    where $\text{rank}_{\text{correct}}(v_{u})$ is the rank position of the correct item $v_{u}$ in the top 10 predictions for user $u$.
D. Implementation Details
In this section, we present the details about the implementation of our SIGMA framework. The corresponding code can be found in the Supplementary Materials. The Mamba block is one of the most important components of the SSM-based model (including SIGMA). So for fair comparison, we set the SSM state expansion factor to 32, the local convolution width to 4, and the block expansion factor to 2 for all models that include Mamba blocks. For the number of stacked layers, we set it to the defaulted 2 for all the selected RNN-based, transformer-based models and Mamba4Rec, comparing them with ECHO and our SIGMA with 1 layer. Moreover, to address the sparsity of Amazon datasets and Yelp dataset, a dropout rate of 0.3 is used, compared to 0.2 for MovieLens-1M.
E. Baselines
(a) GRU4Rec (Vaswani et al. 2017): GRU4Rec utilizes Gated Recurrent Units (GRUs) to capture sequential dependencies within user interaction data. It is particularly effective for session-based recommendations, allowing the model to focus on the most recent and relevant user interactions to make accurate predictions. The model is known for its efficiency in handling session-based data, particularly in capturing user intent during a session (Keles, Wijewardena, and Hegde 2023). (b) BERT4Rec (Kang and McAuley 2018): BERT4Rec adapts the BERT (Bidirectional Encoder Representations from Transformers) architecture for personalized recommendations. Unlike traditional sequential models, BERT4Rec considers both past and future contexts of user behavior by employing a bidirectional self-attention mechanism, which enhances the representation of user interaction sequences. This allows the model to predict items in a more context-aware manner. (c) SASRec (Qin, Yang, and Zhong 2024): SASRec applies a multi-head self-attention mechanism to capture both long-term and short-term user preferences from interaction sequences. It constructs user representations by focusing on relevant parts of the interaction history, thereby improving the quality of recommendations, especially in scenarios where user preferences vary over time. (d) LinRec (de Souza Pereira Moreira et al. 2021): LinRec simplifies the computational complexity of the traditional transformer models by modifying the dot product in the attention mechanism. This reduction in complexity makes LinRec particularly suitable for large-scale recommendation tasks where efficiency is crucial, without significantly sacrificing performance. (e) FEARec (Kim et al. 2019a): FEARec enhances traditional attention mechanisms by incorporating information from the frequency domain. This hybrid approach allows the model to better capture periodic patterns and long-range dependencies in user interaction sequences, leading to more powerful and accurate recommendations. (f) Mamba4Rec (Jannach and Ludewig 2017): Mamba4Rec leverages Selective State Space Models (SSMs) to address the effectiveness-efficiency trade-off in sequential recommendation. By using SSMs, Mamba4Rec efficiently handles long behavior sequences, capturing complex dependencies while maintaining low computational costs. It outperforms traditional self-attention mechanisms, especially in scenarios involving long user interaction histories. (g) EchoMamba4Rec (12): Building on Mamba4Rec, EchoMamba4Rec introduces a frequency domain filter to remove noise and enhance the signal in sequential data. This bi-directional model processes sequences both forward and backward, providing a more comprehensive understanding of user behavior. The Fourier Transform-based filtering further improves the model’s robustness and accuracy in predicting user preferences.
F. Efficiency Comparison
In this section, We will present and analyze the efficiency comparison on other datasets with our proposed SIGMA. From Table 9, we can see that the SSM-based models (including our SIGMA) show consistent efficiency in the other three datasets (Yelp, Sports, and ML-1M). Although for ML-1M, our SIMGA shows an increase in GPU Memory due to our partial flipping method, which means we need to store and compute a new sequence for the reverse direction, our method still achieves higher efficiency remarkably compared with the transformer-based methods, except for LinRec. The experimental results present the fact that our SIGMA can achieve a better efficiency-effectiveness trade-off.
Table 7: Efficiency Comparison: Inference time (ms) per batch and GPU memory (GB).
Dataset Model Infer. GPU Memory
Yelp SASRec 443ms 9.28G
FEARec 483ms 10.01G
LinRec 353ms 7.46G
Mamba 361ms 7.32G
ECHO 368ms 8.46G
SIGMA 352ms 8.27G
Sports SASRec 398ms 7.27G
FEARec 416ms 7.84G
LinRec 294ms 6.07G
Mamba 293ms 5.89G
ECHO 291ms 5.99G
SIGMA 283ms 6.03G
ML-1M SASRec 87ms 21.51G
FEARec 109ms 21.28G
LinRec 59ms 11.79G
Mamba 55ms 6.89G
ECHO 63ms 9.92G
SIGMA 59ms 9.89G
G. Grouped Users Analysis
In this section, we will present the grouped user analysis on the other three datasets i.e., ML-1M, Yelp and Games. We also group the user by their interaction length, followed by our experiment setting. The detailed distribution is listed in Table 8. The performance of the selected models is presented in Figure 7. Noted that, as illustrated in Table 8, the interaction length of users in ML-1M are all longer than 5, so the performance for the “short” group in ML-1M is reasonably empty. All the experiments above further prove the superiority of our proposed SIGMA framework on all groups, showing the effectiveness of our FE-GRU and PF-Mamba in dealing with context modeling and short sequence modeling, respectively.
Figure 7: Grouped Users Analysis on ML-1M, Yelp and Games.
Table 8: User sample distribution
Dataset Short(0-5) Medium(5-20) Long(20-inf)
Games 27631 25084 2426
Yelp 34659 42934 5306
ML-1M 0 177 5863
H. Ablation Study
In this section, we will analyze the efficacy of three key components with SIGMA on other datasets i.e., Sports, Games, Yelp and ML-1M. According to the data statistics in the experiment setting, we can clearly see that Yelp, Sports, Beauty, and Games have similar average interaction lengths. Correspondingly, the ablation study on Sports, Games, and Yelp datasets shows a similar tendency to the one on Beauty. But for ML-1M, due to the relatively long sequence, we can find that FE-GRU contributes the least since it mainly focuses on enhancing the hidden representation for short sequences, while the other two components show a remarkable drop when removing them, proving the effectiveness of our designed PF-Mamba in contextual information modeling.
Table 9: Ablation study on other datasets.
Dataset Methods HR@10 NDCG@10 MRR@10
Sports Default 0.0735 0.0590 0.0556
w/o partial flipping 0.0711 0.0582 0.0544
w/o DS gate 0.0713 0.0577 0.0529
w/o FE-GRU 0.0622 0.0481 0.0473
Games Default 0.1627 0.1088 0.0924
w/o partial flipping 0.1601 0.1067 0.0911
w/o DS gate 0.1597 0.1054 0.0889
w/o FE-GRU 0.1562 0.0978 0.0861
Yelp Default 0.0629 0.0412 0.0346
w/o partial flipping 0.0611 0.0402 0.0327
w/o DS gate 0.0613 0.0389 0.0314
w/o FE-GRU 0.0595 0.0377 0.0301
ML-1M Default 0.3308 0.1906 0.1479
w/o partial flipping 0.3289 0.1887 0.1459
w/o DS gate 0.3288 0.1866 0.1443
w/o FE-GRU 0.3304 0.1901 0.1471
I. Guideline for reproducibility
In this section, we will provide detailed guidelines for reproducing our results. In the model file we released, we have baseline and baseline_config folders, which respectively contain almost all the baseline methods (others can be easily found in recbole’s original models) and corresponding configurations mentioned in the overall experiment; datasets folder storing the chosen datasets (Beauty, Sports, Games, Yelp and ML-1M) in atomic files to fit the recbole framework (Wang, He, and Zhu 2024); model file containing the gated_mamba.py, which is the main structure of our SIGMA. To reproduce our proposed SIGMA framework, you can directly run the run.py file with the proper command. It is noteworthy that the environment and version of mamba_ssm are quite important in reproducing the experimental results, so please check your environment referencing the requirements.yaml file. Specifically, you are recommended to run run.py for successful reproduction, which contains the calls to gated_mamba.py. For the relevant parameters to perform the experiments in hyperparameter analysis, you can directly set it in config.yaml with other dataset settings and model settings. We also attach the RecMamba.ipynb to show the raw training procedure in Colab.